OpenStudy (anonymous):

Please help I still don't understand this Algebra 1 question please help. Solve the equation 3x-7=5x+19

OpenStudy (anonymous):

Alrighty lets see... 3x-7=5x+19 we want to try to move things around so we get x= a number so then 3x-7=5x+19 add 7 3x=5x+26 now subtract 5x -2x=26 you got it from there? finish solving for x :)

OpenStudy (anonymous):

I think so if not I will ask for help again

OpenStudy (anonymous):

Just so you can check your work: -2x=26 divinde by -2 x= answer :)

OpenStudy (anonymous):

3x-7=5x+19 3x-5x=19+7 -2x=26 X=26 divided by -2 X=-13